# Reconstruction of Caterpillar Tanglegrams

### v1

@@ 4. Reconstruction of caterpillar tanglegrams
    Finally, assume that exactly one of $v_{n}$ and $v_{n-1}$ is matched to $Q$. Without loss of generality assume that $v_{n}$ is matched to $Q$ and $v_{n-1}$ is matched to a strippable leaf of $R$. Now notice that if $v_{n-2}$ is matched to $Q$, then there is precisely one card with $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ both matched to $Q$. If $v_{n-2}$ is matched to a strippable leaf of $R$, then there is no such card. Thus, we can distinguish these two subcases using $\mathcal{D}^{m}(\mathcal{T})$.
    1. (a)
      If $v_{n-2}$ is matched to $Q$, there is a unique card, $\mathcal{T}-v_{n-1}$, with the property that both of $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to $Q$. Furthermore, this card has the property that all leaves of $Q$ are present and we can use it to determine the matching of all leaves of $C_{n}$ that are matched to $Q$. Now, any other card with the property that $R^{\prime}$ has at least $i$ strippable leaves (meaning that a leaf from $Q$ was removed) allows us to determine the matching of all leaves of $C_{n}$ matched to the strippable leaves of $R$. Combining these cards, we can reconstruct $\mathcal{T}$.
    2. (b)
      If $v_{n-2}$ is matched to a strippable leaf of $R$, let $v_{n-1},v_{n-2},\ldots,v_{n-t}$ with $12\leq t\leq i+$ 1 be a maximal sequence of $t$ leaves all matched to strippable leaves of $R$, whereas $v_{n-t-1}$ is matched to $Q$. Using similar reasoning as in the previous cases, we can determine $t$.
      Now, since $|Q|\geq 4$, we can find a card such that $R^{\prime}$ has at least $i$ strippable leaves (implying that a leaf from $Q$ has been removed), $v^{\prime}_{n-1}$ is matched to $Q$, and there is a maximal sequence of at least $t$ leaves $v^{\prime}_{n-2},v^{\prime}_{n-3},\ldots,v^{\prime}_{n-t-1},\ldots$ all matched to strippable leaves. We can use this card to determine the indices (in the distance-to-root labeling) and matching of the leaves of $C_{n}$ matched to the strippable leaves of $R$. Furthermore, we can identify a second card with $v_{n-1}^{\prime}$ matched to $Q$ and a maximal sequence of only $t-1$ leaves $v_{n-2}^{\prime},\ldots,v^{\prime}_{n-t}$ matched to strippable leaves of $R$ (implying that a strippable leaf has been removed while all leaves of $Q$ are present). This card allows us to determine the matching for all remaining leaves and we can reconstruct $\mathcal{T}$.
We next consider the case that $R$ is type 2.
@@
Let $\mathcal{T}=(C_{n},R,\sigma)$ be a tanglegram of size $n\geq 5$ such that $R$ is type 2. Then, $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(C_{n},R,\sigma)$.
@@
Let $\{v_{1},\ldots,v_{n}\}$ be a distance-to-root labeling of the leaves of $C_{n}$ and, on each card, let $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ be a distance-to-root labeling of the leaves of $C_{n-1}$. Let $R=R_{1}\oplus R_{2}$ with $|R_{1}|,|R_{2}|\geq 2$ (since $R$ is type 2). As $n\geq 5$, we can reconstruct $C_{n}$ and $R$ from $\mathcal{D}^{m}(\mathcal{T})$ (Theorem 2.7).
We first argue that we can determine from $\mathcal{D}^{m}(\mathcal{T})$ whether $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree $R_{i}$ of $R$ or not. If $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree of $R$, then on at least $n-2\geq 3$ cards, leaves $v^{\prime}_{n-2}$ and $v^{\prime}_{n-1}$ are matched to the same maximal pending subtree of $R^{\prime}$. If $v_{n}$ and $v_{n-1}$ are not matched to the same maximal pending subtree of $R$, then on precisely one card, leaves $v^{\prime}_{n-2}$ and $v^{\prime}_{n-1}$ are matched to the same maximal pending subtree of $R^{\prime}$. Thus, we can distinguish these two cases.
@@
    First, assume that $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree of $R$. Let $v_{n},v_{n-1},\ldots,v_{n-s}$ with $31\leq s\leq n-$ 3 be a maximal sequence of $s+1$ leaves all matched to the same maximal pending subtree of $R$, say $R_{i}$, whereas $v_{n-s-1},\ldots,v_{n-t}$ with $s+1\leq t\leq n-1$ is a maximal sequence of $t-s\geq 1$ leaves matched to $R_{3-i}$.
    In this case, there are precisely $s+1$ cards (corresponding to the removal of $v_{n},v_{n-1},\ldots,v_{n-s}$, respectively) with the property that there is a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ all matched to $R_{i}$, whereas $v^{\prime}_{n-s-1},\ldots,v^{\prime}_{n-t}$ is a maximal sequence of $t-s$ leaves matched to $R_{3-i}$ (on the remaining $n-s-1\geq 1$ cards, at least $s+1$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ are matched to $R_{i}$). Thus, we can determine $s$ and $t$.
    Now, using a card containing a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ matched to $R_{i}$ (meaning that a leaf from $R_{i}$ was removed but all leaves from $R_{3-i}$ are present), allows us to determine which of the leaves $v_{1},\ldots,v_{n-s-1}$ are matched to $R_{3-i}$ and which are matched to $R_{i}$. Furthermore, for those matched to $R_{3-i}$, we can determine the matching.
    A second card with a maximal sequence of at least $s+1$ leaves
    $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ all matched to $R_{i}$ and a maximal sequence of $t-s-1$ of the next (lower-indexed) leaves matched to $R_{3-i}$ (meaning that a leaf from $R_{3-i}$ was removed but all leaves from $R_{i}$ are present), allows us to determine the matching for all remaining leaves. Thus, we can reconstruct $\mathcal{T}$ in this case.
@@
    Second, assume that $v_{n}$ and $v_{n-1}$ are matched to different maximal pending subtrees of $R$. Without loss of generality assume that $v_{n}$ is matched to $R_{i}$ and $v_{n-1},v_{n-2}$ are matched to $R_{3-i}$. In particular, let $v_{n-1},v_{n-2},\ldots,v_{n-t}$ with $22\leq t\leq n-$ 2 be a maximal sequence of $t$ leaves all matched to $R_{3-i}$, whereas $v_{n-t-1}$ is matched to $R_{i}$. Using similar reasoning as in the previous case, we can determine $t$.
    Now, there is a unique card, $\mathcal{T}-v_{n}$, with a maximal sequence of $t$ leaves $v^{\prime}_{n-1},v^{\prime}_{n-2},\ldots,v^{\prime}_{n-t}$ matched to $R_{3-i}$ and we can use it to determine the indices (in the distance-to-root labeling) and matching of the leaves of $C_{n}$ matched to leaves from $R_{3-i}$, except for $v_{n-1}$ and $v_{n-2}$. Let $r$ denote the size of the subtree this sequence of $t$ leaves is matched to on this card. To resolve the matching for $v_{n-1}$ and $v_{n-2}$, we identify a second card with the property that $v^{\prime}_{n-1}$ is matched to $R_{i}$, there is a maximal sequence of $t$ leaves $v^{\prime}_{n-2},v^{\prime}_{n-3},\ldots,v^{\prime}_{n-t-1}$ matched to $R_{3-i}$, and $|R_{3-i}|=r$ (implying that a leaf from $R_{i}$ not matched to $v_{n}$ was removed; such a card must exist since $|R_{i}|\geq 2$). This card allows us to determine the matching of $v_{n-1}$ and $v_{n-2}$. Finally, we can identify a third card with $v_{n-1}^{\prime}$ matched to $R_{i}$ and a maximal sequence of only $t-1$ leaves $v_{n-2}^{\prime},\ldots,v^{\prime}_{n-t}$ matched to $R_{3-i}$ (implying that a leaf from $R_{3-i}$ was removed). This card allows us to determine the matching for all remaining leaves and we can reconstruct $\mathcal{T}$.

### v2

@@ 7. Caterpillar tanglegrams with type 2 left or right tree
    $R$ is type 2c or $R$ is type 2b and the two subtrees of $R^{\prime}$ are of different sizes:
    We can identify $R^{\prime}_{1},R^{\prime}_{2}$ correctly on this card.
    Choose $j$ such that $v^{\prime}_{n},v^{\prime}_{n-1}$ are matched to the same subtree $R^{\prime}_{j}$ of $R^{\prime}$ in the card. If $v_{n},v_{n-1}$ are matched to the same subtree of $R$, then $i=j$. Otherwise i=3ji=3-j.
  2. (2)
    $R$ is type 2b and the two subtrees of $R^{\prime}$ are of the same size:
    In this case the card was obtained by removing a leaf from $R_{2}$.
    If $v_{n},v_{n-1}$ are matched to the same pending subtree of $R$, then the maximality condition ensures this card contains all leaves of this pending subtree and $i=1$.
    If $v_{n},v_{n-1}$ are matched to different subtrees of $R$, then this card is $\mathcal{T}-v_{n}$, so $i=2$.
Thus, we can determine $i$ from $\mathcal{D}^{m}(\mathcal{T})$.
@@
Let $\mathcal{T}=(C_{n},R,\sigma)$ be a tanglegram of size $n\geq 6$ such that $R$ is type 2. If the leaves of the cherry of the left tree are matched to the same pending subtree of $R$, then $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
@@
As $n\geq 6$, we can reconstruct $C_{n}$ and $R$ from $\mathcal{D}(\mathcal{T})$ (Lemma 2.17). By Lemma  7.1, $\mathcal{D}^{m}(\mathcal{T})$ determines that $v_{n},v_{n-1}$ are matched to the same subtree $R_{i}$ of $R$, and also the index $i$ of this subtree. Therefore the class of tanglegrams of size at least $6$ with left tree caterpillar and right tree type 2 tree and the leaves of the cherry are matched to the same pending subtree of the right tree is a deck-decidable class.
In addition, in any card in which the leaves $v^{\prime}_{n-1},v^{\prime}_{n-2}$ of the left tree $C_{n-1}$ are matched to the same subtree of $R^{\prime}$, we can determine the indices $R^{\prime}_{1}$, $R^{\prime}_{2}$ correctly, as we know the value of $i$ and $R^{\prime}_{i}$ is the subtree of $R^{\prime}$ that both $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to.
Choose $s,t$ such that the $s+1$ leaves in the sequence $v_{n},v_{n-1},\ldots,v_{n-s}$ are all matched to the same pending subtree $R_{i}$ of $R$, the tst-s leaves in the sequence $v_{n-s-1},\ldots,v_{n-t}$ are matched to $R_{3-i}$, and either t=n1t=n-1 or $v_{n-t-1}$ is matched to $R_{i}$. We have $1\leq s\leq n-3$ and $s+1\leq t\leq n-1$.
We start with showing that we can determine $s,t$ from $\mathcal{D}^{m}(\mathcal{T})$.
If $s=1$, then in exactly two cards ($\mathcal{T}-v_{n}$ and $\mathcal{T}-v_{n-1}$), $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to different pending subtrees of $R^{\prime}$. If $s\geq 2$, then in all cards, we have $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ matched to the same pending subtree of $R^{\prime}$. Moreover, when $s\geq 2$, in exactly $s+1$ cards (of the form $\mathcal{T}-v_{j}$ where $n-s\leq j\leq n$), $v^{\prime}_{n-1}$ and $v^{\prime}_{n-s-1}$ are not matched to the same pending subtree of $R^{\prime}$. Thus, $\mathcal{D}^{m}(\mathcal{T})$ determines the value of $s$. Moreover, if $s\geq 2$, we can identify $R^{\prime}_{1}$ and $R^{\prime}_{2}$ in every card, and if $s=1$, we can identify $R^{\prime}_{1}$ and $R^{\prime}_{2}$ in every card different from $\mathcal{T}-v_{n}$ and $\mathcal{T}-v_{n-1}$.
If $s\geq 2$, then in all of the $s+1$ cards where $v^{\prime}_{n-1}$ and $v^{\prime}_{n-s-1}$ are matched to different subtrees of $R^{\prime}$, every leaf in the sequence $v^{\prime}_{n-s-1},\ldots,v^{\prime}_{n-t}$ is matched to the same pending subtree of $R^{\prime}$ while either t=n1t=n-1 or $v^{\prime}_{n-t-1}$ is matched to a different pending subtree, so we can determine $t$ from $\mathcal{D}^{m}(\mathcal{T})$.
If $s=1$ and $t=1$, then there is precisely one card ($\mathcal{T}-v_{n-2}$) in which $v^{\prime}_{n-1},v^{\prime}_{n-2}$, and $v^{\prime}_{n-3}$ are all matched to the same pending subtree of $R^{\prime}$. If $s=1$ and $t\geq 2$, then in none of the cards we have $v^{\prime}_{n-1},v^{\prime}_{n-2}$, and $v^{\prime}_{n-3}$ all matched to the same pending subtree of $R^{\prime}$.
If $s=1$ and $t\geq 2$, there are precisely two cards $(\mathcal{T}-v_{n}$ and $\mathcal{T}-v_{n-1}$) such that $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are not matched to the same pending subtree of $R^{\prime}$. Consider these two cards only. If t=n1t=n-1, then $v^{\prime}_{1},v^{\prime}_{2},\ldots,v^{\prime}_{n-3}$ are all matched to the same pending subtree of $R^{\prime}$. If t<n1t<n-1, then $v^{\prime}_{n-3}$ and $v^{\prime}_{n-t-1}$ are matched to different pending subtrees of $R^{\prime}$, but for each $j$ with $3\leq j\leq t$, $v^{\prime}_{n-3}$ and $v^{\prime}_{n-j}$ are matched to the same pending subtree of $R^{\prime}$. Therefore, when $s=1$, $\mathcal{D}^{m}(\mathcal{T})$ determines the value of $t$.
Moreover, if $s=1$, then in the cards where $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are not matched to the same subtree of $R^{\prime}$ (which are $\mathcal{T}-v_{n}$ and $\mathcal{T}-v_{n-1}$), leaf $v^{\prime}_{n-3}$ is matched to $R^{\prime}_{i}$ when $t=1$ and it is matched to $R^{\prime}_{3-i}$ when $t\geq 2$. Therefore, even when $s=1$, we can identify the proper indices for $R^{\prime}_{1}$ and $R^{\prime}_{2}$ in all cards. Without loss of generality, when $s=1$ and $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to different subtrees, we choose the distance-to-root labeling of the left tree $C_{n-1}$ such that $v^{\prime}_{n-1}$ is matched to $R_{i}$, so $v^{\prime}_{n-2}=v^{\prime}_{n-s-1}$ is matched to $R_{3-i}$.
Now choose a card in which $v^{\prime}_{n-s-1}$ is matched to $R_{3-i}$. This means that a leaf in $R_{i}$ that was matched to one of the $s+1$ leaves $v_{n},v_{n-1},\ldots,v_{n-s}$ was removed but all leaves from $R_{3-i}$ are present. This allows us to determine both $X_{1}$ and $X_{2}$, and, as $v_{n},v_{n-1}\notin X_{3-i}$, this card also determines the matching between $X_{3-i}$ and $R_{3-i}$.
Next choose a card in which $v^{\prime}_{n-s-1}$ is matched to $R_{i}$, implying that a leaf from $R_{3-i}$ was removed to obtain $R^{\prime}_{3-i}$ (as we know $R_{1},R_{2}$ and $i$, and can identify $R^{\prime}_{1},R^{\prime}_{2}$ correctly, such a card can be found). Since we know the set $X_{i}$, and $v_{n},v_{n-1}\in X_{i}$, this card allows us to determine the matching between $X_{i}$ and $R_{i}$, while for the two vertices matched to $v_{n}$ and $v_{n-1}$, we may not know which vertex is matched to $v_{n}$ and which to $v_{n-1}$. However, the two possible matchings give the same unlabeled tanglegram. Therefore, $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$. ∎
Lemma 7.3.
Let $\mathcal{T}=(C_{n},R,\sigma)$ be a tanglegram of size $n\geq 6$. If $R$ is type 2, $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
Proof.
As $n\geq 6$, we can reconstruct $C_{n}$ and $R$ from $\mathcal{D}(\mathcal{T})$ by Lemma 2.17; and the class of tanglegrams of size at least $6$ where the left tree is a caterpillar and the right tree is type 2 is deck-decidable. Lemma 7.1 gives that $\mathcal{D}^{m}(\mathcal{T})$ determines whether $v_{n},v_{n-1}$ are matched to the same subtree of $R$, and also the index $i$. If $v_{n},v_{n-1}$ are matched to the same subtree of $R$, we are done by Lemma 7.2.
Assume that $v_{n}$ and $v_{n-1}$ are matched to different pending subtrees of $R$. Recall that we choose a distance-to-root labeling such that if $v_{n}$ and $v_{n-1}$ are matched to different subtrees of $R$, then $v_{n-2},v_{n-1}$ are matched to the same subtree $R_{3-i}$. Let $t$ be such that the $t$ leaves in the sequence $v_{n-1},v_{n-2},\ldots,v_{n-t}$ are all matched to $R_{3-i}$ while $v_{n-t-1}$ is matched to $R_{i}$. We have $2\leq t\leq n-2$.
We have exactly one card (namely $\mathcal{T}-v_{n}$) in which $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to the same pending subtree of $R^{\prime}$. Consider this card: all leaves in the sequence $v^{\prime}_{n-1},v^{\prime}_{n-2},\ldots,v^{\prime}_{n-t}$ are matched to $R^{\prime}_{3-i}$ and $v^{\prime}_{n-t-1}$ is matched to $R^{\prime}_{i}$. In any other card, $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to different pending subtrees. Therefore, we can determine $t$ from $\mathcal{D}^{m}(\mathcal{T})$.
Take the unique card in which $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to the same subtree $R^{\prime}_{3-i}$. As this card is $\mathcal{T}-v_{n}$, $v_{n}\in X_{i}$, and $v_{n-1}\in X_{3-i}$, we can determine $X_{1}$ and $X_{2}$, and the pending subtree $R_{3-i}$ of $R$ (uniquely, unless $R$ is type 2a).
Assume first that $t\geq 3$. Consider only the n1n-1 cards in which $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to different subtrees of $R^{\prime}$, and without loss of generality, use the labeling in which $v^{\prime}_{n-2}$ and $v^{\prime}_{n-3}$ are matched to the same subtree of $R^{\prime}$. As $t\geq 3$, this ensures that $v^{\prime}_{n-1}$ corresponds to the vertex $v_{n}$ and $v^{\prime}_{n-2},v^{\prime}_{n-3}$ are matched to the pending subtree $R^{\prime}_{3-i}$. This ensures that the pending subtrees $R^{\prime}_{1}$, $R^{\prime}_{2}$ are correctly identified in these cards.
Find a card in which the size of the subtree $v^{\prime}_{n-2}$ is matched to is $\left\langle R_{3-i}\right\rangle=|X_{3-i}|$. As $v_{n}\notin X_{3-i}$, and this card is of the form $\mathcal{T}-v_{j}$ for some $v_{j}\in X_{i}-\{v_{n}\}$, this card allows us to determine the matching between vertices of $X_{3-i}$ and $R_{3-i}$.
Now find a card in which the size of the subtree $v^{\prime}_{n-1}$ is matched to is $\left\langle R_{i}\right\rangle=|X_{i}|$. As $v_{n-1},v_{n-2}\notin X_{i}$, and this card is of the form $\mathcal{T}-v_{j}$ for some $v_{j}\in X_{3-i}$, this card allows us to determine the matching between vertices of $X_{i}$ and $R_{i}$. Thus, we can reconstruct $\mathcal{T}$ when $t\geq 3$.
Assume in what remains that $t=2$. We have the following cases:
@@
    $R$ is type 2c:
    Since we know $R_{1},R_{2}$, and $i$, and we can correctly identify $R^{\prime}_{1}$ and $R^{\prime}_{2}$ in every card, in every card different from $\mathcal{T}-v_{n}$, we can choose the distance-to-root labeling such that $v^{\prime}_{n-1}$ is matched to $R^{\prime}_{i}$ and $v^{\prime}_{n-2}$ is matched to $R^{\prime}_{3-i}$. Consider the cards different from $\mathcal{T}-v_{n}$ and take one in which a leaf from $R_{3-i}$ was removed. As in this card, $v^{\prime}_{n-1}$ is matched to $R_{i}$ and $v^{\prime}_{n-2}$ is matched to $R^{\prime}_{3-i}$, this gives us the matching between $X_{i}$ and $R_{i}$. Take another card that is not $\mathcal{T}-v_{n}$, in which a leaf from $R_{i}$ was removed — this gives us the matching between $X_{3-i}$ and $R_{3-i}$. This determines $\mathcal{T}$.
@@
    $R$ is type 2b:
    We know $i$, $R_{1},R_{2}$, and we can identify $R_{1}^{\prime}$, $R_{2}^{\prime}$ correctly in every card in which $R^{\prime}$ is not type 2a, Moreover, $\left\langle R_{1}\right\rangle\geq 3$, $\left\langle R_{2}\right\rangle=\left\langle R_{1}\right\rangle+1\geq 4$, and if $R^{\prime}$ is type 2a, then a leaf from $R_{2}$ was removed to obtain $R^{\prime}$.
    Take a card from which a leaf from $R_{1}$ was removed (so $R^{\prime}_{1}$ and $R^{\prime}_{2}$ can be identified) and $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to different pending subtrees of $R^{\prime}$. This card determines the matching between $X_{2}$ and $R_{2}$.
    Now consider $Z=X_{2}-\{v_{1},v_{n},v_{n-2},v_{n-1}\}$. As $|X_{2}|=\left\langle R_{2}\right\rangle\geq 4$ and only one of $v_{n},v_{n-1}$ is in $X_{2}$, $Z\neq\emptyset$. For any $v_{j}\in Z$, card $\mathcal{T}-v_{j}$ has $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ matched to different pending subtrees of $R^{\prime}$. Without loss of generality, choose the labeling such that $v^{\prime}_{n-2}$ and $v^{\prime}_{n-3}$ are matched to the same pending subtree of $R^{\prime}$. Then $v^{\prime}_{n-1}$ and $v^{\prime}_{1}$ are matched to the same pending subtree of $R^{\prime}$ precisely when $v_{1}\in X_{i}$.
    Consider any card where a leaf from $R_{2}$ has been removed, $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to different pending subtrees of $R^{\prime}$, and $v^{\prime}_{1}$ is matched to the opposite subtree as $v^{\prime}_{n-3}$ exactly if $v_{1}\in X_{i}$. We may choose the labeling so that $v^{\prime}_{n-2}$ and $v^{\prime}_{n-3}$ are matched to the same pending subtree and so $v^{\prime}_{1}$ and $v^{\prime}_{n-1}$ are matched to the same subtree exactly when $v_{1}\in X_{i}$. This card is of the form $\mathcal{T}-v_{p}$ for some $p\neq n$.
    We will first show that if $v_{p}\notin Z$, then $p=1$ and $v_{1},v_{2}$ are matched to the same subtree of $R$.
    If $v_{p}\notin Z$, then $p\in\{n-1,n-2,1\}$.
    Assume $p\in\{n-1,n-2\}$. Since $t=2$, we have that $v^{\prime}_{n-3}$ corresponds to $v_{n-3}$, $v^{\prime}_{n-1}$ corresponds to one of $v_{n-1},v_{n-2}$ (the other is $v_{p}$), $v^{\prime}_{n-2}$ corresponds to $v_{n}$, and $v^{\prime}_{1}$ corresponds to $v_{1}$. However, this contradicts the fact that $v^{\prime}_{n-1}$ and $v^{\prime}_{1}$ match to the same subtree of $R^{\prime}$ precisely when $v_{1}\in X_{i}$. Therefore, if $v_{p}\notin Z$, then $p=1$.
    Assume $p=1$. The card we choose is $\mathcal{T}-v_{1}$, so for each $j\in\{1,\ldots,n-1\}$, $v^{\prime}_{n-j}$ corresponds to $v_{n+1-j}$, and $v_{n},v_{1}$ are matched to the same subtree of $R$ precisely when $v_{n},v_{2}$ are matched to the same subtree of $R$. Therefore, $v_{1},v_{2}$ are matched to the same subtree of $R$.
    Thus, whether or not $p\in Z$, this card allows us to determine the matching between $R_{1}$ and $X_{1}$. Thus, $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
  3. (3)
    $R$ is type 2a:
    In this case $i=1$, $n$ is even, $|X_{1}|=|X_{2}|=\left\langle R_{1}\right\rangle=\left\langle R_{2}\right\rangle\geq 3$, $v_{n},v_{n-3}\in X_{1}$ and $v_{n-1},v_{n-2}\in X_{2}$. Note that for $j\in\{1,2\}$, a matching between $X_{j}$ and $R_{j}$ is an ordering $w_{1},\ldots,w_{\frac{n}{2}}$ of the leaves of $R_{j}$ such that for each $q\in\left\{1,\ldots,\frac{n}{2}\right\}$, we have that $v_{k_{q,j}}$ is matched to $w_{q}$.
    Consider the n1n-1 cards in which $v^{\prime}_{n-1},v^{\prime}_{n-2}$ are matched to different subtrees of $R^{\prime}$ (so $v_{n}$ was not removed). Each of these give an ordering of the leaves of the larger pending subtree of $R^{\prime}$, so one of the two possible matchings from $X_{1}$ to $R_{1}$ or $X_{2}$ to $R_{2}$.
    If all n1n-1 of these orderings are the same (up to isomorphism of the subtree), $\mathcal{T}$ is determined.
    Otherwise, one ordering appears $\frac{n}{2}-1$ times, and the other $\frac{n}{2}$ times. As $i=1$, $v_{n}\in X_{1}$, and $\mathcal{T}-v_{n}$ is not among the n1n-1 cards considered, the ordering that appears fewer times gives the matching between $X_{2}$ and $R_{2}$. $\mathcal{T}$ is determined.
8. Conclusion